OpenStudy (anonymous):

simplify sin2x/2cosx with steps

OpenStudy (solomonzelman):

\[\frac{Sin^2x}{2Cosx}\]\[\color{red} {Tanx=\frac{Sinx}{Cosx}}\]

OpenStudy (solomonzelman):

use the definition in red.

OpenStudy (loser66):

hold on, clarify sin^2 x or sin (2x)

OpenStudy (anonymous):

i bet $7 it is \(\sin(2x)\)

OpenStudy (anonymous):

it is sin (2x)

OpenStudy (anonymous):

i win

OpenStudy (solomonzelman):

\[\frac{Sinx \times Sinx}{Cosx}~~~~>~~~~\frac{Sinx}{Cosx}\times Sinx\]

OpenStudy (loser66):

ahhahah @satellite73 medal for your bet

OpenStudy (anonymous):

\[\sin92x=2\sin(x)\cos(x)\] then a bit of cancellation will finish

OpenStudy (anonymous):

typo there \[\sin(2x)=2\sin(x)\cos(x)\]

OpenStudy (solomonzelman):

I am confused, what is the question?

OpenStudy (loser66):

the question is simplify , but you started the wrong problem

OpenStudy (solomonzelman):

what is the problem, (the equation to simplify) ?

OpenStudy (anonymous):

its simplify sin(2x)/2cosx, i am unsure of what to do with the 2cosx

OpenStudy (loser66):